# On Hardy-Littlewood-Sobolev estimates for degenerate Laplacians

### v1

@@ 1. Introduction
Degenerate elliptic operators where the degeneracy is controlled by weights in the Muckenhoupt class have been introduced in [FKS82]. Since then, they have attracted attention with extension to parabolic equations [CS84a, CS84b, CS85, Ish99]. For $A_{2}(\mathbb{R}^{n})$ weights $\omega$, $-\Delta_{\omega}$ can be defined as a maximal accretive operator on the weighted space $L^{2}_{\omega}(\mathbb{R}^{n})$. Much is known. For example, the associated semigroup is bounded on $L^{p}_{\omega}(\mathbb{R}^{n})$ for all $1\leq p\leq\infty$, it has $H^{\infty}$ functional calculus (equivalently, Borel functional calculus) if $1<p<\infty$. The particular study of its square root can be found in [CUR15, CUMR18] with a proof of the Kato problem on $L^{2}_{\omega}(\mathbb{R}^{n})$ and extension to some range of $p$, and even a "weighted theory" with weights in the Muckenhoupt classes for the measure $w(x)\,dx$.
@@
Inequalities (1.1) and (1.2) are a priori weighted Sobolev embeddings of a functional nature. Some more geometric weighted Sobolev-Gagliardo-Nirenberg inequalities were obtained in [LMPT10]. Namely, for all $1\leq p<q<\infty$ with $\frac{1}{2}=\frac{n}{2}(\frac{1}{p}-\frac{1}{q})$, $\|f\|_{L^{q}_{{\tilde{\omega}}^{q}}(\mathbb{R}^{n})}\lesssim\|\nabla f\|_{L^{p}_{\tilde{\omega}^{p}}(\mathbb{R}^{n})}$ when $\tilde{\omega}\in A_{p,q}(\mathbb{R}^{n})$ (which, again, requires $\tilde{\omega}^{p}\in RH_{\frac{q}{p}}(\mathbb{R}^{n})$).
We can compare these results with ours only when $\alpha=1/2$, for $p=2$ where $\|(-\Delta_{\omega})^{1/2}f\|_{L^{2}_{\omega}(\mathbb{R}^{n})}=\|\nabla f\|_{L^{2}_{\omega}(\mathbb{R}^{n})}$ by construction and in the ranges of $p$ where $\|\nabla f\|_{L^{p}_{\omega}(\mathbb{R}^{n})}$ dominates or is dominated by $\|(-\Delta_{\omega})^{1/2}f\|_{L^{p}_{\omega}(\mathbb{R}^{n})}$. Deducing our results from them or vice-versa also require some further restrictions on the class of weights.
Coming back to our result, we will show that the reverse Hölder assumption is sharp in the sense that it is implied by (1.1). At the level of the semigroup, we shall see that with the assumptions of the theorem, then $\omega\in RH_{\frac{q}{p}}(\mathbb{R}^{n})$ is equivalent to $e^{t\Delta_{w}}:L^{p}_{\omega}(\mathbb{R}^{n})\to L^{q}_{{\omega}^{q/p}}(\mathbb{R}^{n})$ with bound on the order $t^{-\alpha}$ for all $t>0$. This polynomial bound is the key point of our method.
[…]
@@
As a matter of motivation for this work, bounds for fractional weighted Laplacians are needed to study second order degenerate parabolic equations with unbounded lower order terms. The second author will do it in a forthcoming work.
@@
Throughout this paper, we fix an integer $n\geq 1$. We use the notation $\mathrm{d}x$ for the Lebesgue measure on $\mathbb{R}^{n}$. For any Lebesgue measurable set $E\subset\mathbb{R}^{n}$, we denote its Lebesgue measure by $|E|$. By convention, the notation $C(a,b,\ldots)$ denotes a constant depending only on the parameters $(a,b,\ldots)$. All cubes $Q\subset\mathbb{R}^{n}$ considered in this paper are assumed to have sides parallel to the coordinate axes.
@@ Acknowledgements
We want to thank Chema Martell for bringing our attention to [CMM22] concerning Lorentz spaces after we completed this work. We would also like to thank Carlos Pérez for enlightening discussions.
@@ 2.1. Muckenhoupt weights
(2.1) $[\omega]_{A_{p}}:=\sup_{Q\subset\mathbb{R}^{n}}\left(\frac{1}{|Q|}\int_{Q}\omega(x)\ \mathrm{d}x\right)\left(\frac{1}{|Q|}\int_{Q}\omega^{1-p^{\prime}}(x)\ \mathrm{d}x\right)^{p-1}<\infty,$
@@
$\left(\frac{1}{|Q|}\int_{Q}\omega^{q}(x)\ \mathrm{d}x\right)^{\frac{1}{q}}\leq C\frac{1}{|Q|}\int_{Q}\omega(x)\ \mathrm{d}x,$
for some constant $C<\infty$ and for all cubes $Q\subset\mathbb{R}^{n}$. The infimum of such constants $C$ is what we denote by $[\omega]_{RH_{q}}$. For $q=\infty$, the reverse Hölder class $RH_{\infty}(\mathbb{R}^{n})$ is defined as the set of all weights $\omega$ verifying
$\omega(y)\leq C\frac{1}{|Q|}\int_{Q}\omega(x)\ \mathrm{d}x,\quad\text{for almost every}\ y\in Q,$
@@
$[\omega]_{A_{p,q}}:=\sup_{Q\subset\mathbb{R}^{n}}\left(\frac{1}{|Q|}\int_{Q}\omega^{q}(x)\ \mathrm{d}x\right)^{\frac{1}{q}}\left(\frac{1}{|Q|}\int_{Q}\omega^{-p^{\prime}}(x)\ \mathrm{d}x\right)^{\frac{1}{p^{\prime}}}<\infty,$
@@
It is well known (see, for instance, [Gra08, Example 7.1.7]) that a weight of the form $\omega(x)=|x|^{\beta}$ belongs to the Muckenhoupt class $A_{p}(\mathbb{R}^{n})$ for $p>1$ if and only if $-n<\beta<n(p-1)$. Furthermore, according to point (6) of Proposition 2.1 above, a weight $\omega$ belongs to the intersection $A_{2}(\mathbb{R}^{n})\cap RH_{\frac{q}{p}}(\mathbb{R}^{n})$ for $p<q$ if and only if $\omega^{\frac{q}{p}}\in A_{1+\frac{q}{p}}(\mathbb{R}^{n})$. As a result, a typical example of a weight $\omega$ considered in the statements of the results in this paper, satisfying $\omega\in A_{2}(\mathbb{R}^{n})\cap RH_{\frac{q}{p}}(\mathbb{R}^{n})$ with $p<q$, is given by
@@ 2.2. Function spaces
During all this paper, $\omega$ denotes a fixed weight belonging to the Muckenhoupt class $A_{2}(\mathbb{R}^{n})$. By definition (2.1) with $p=2$, the weight $\omega^{-1}$ is also in the Muckenhoupt class $A_{2}(\mathbb{R}^{n})$ with $[\omega^{-1}]_{A_{2}}=[\omega]_{A_{2}}$. We introduce the measure $\mathrm{d}\omega:=\omega(x)\mathrm{d}x$ and if $E\subset\mathbb{R}^{n}$ a Lebesgue measurable set, we write $\omega(E)$ instead of $\int_{E}\mathrm{d}\omega$. For all $x\in\mathbb{R}^{n}$ and $r>0$, we set $\omega_{r}(x):=\omega(B(x,\sqrt{r}))$.
It follows from (2.1) with $p=2$ that there exits constants $\eta\in(0,1)$ and $\beta>0$, depending only on $n$ and $[\omega]_{A_{2}}$, such that
@@
whenever $Q\subset\mathbb{R}^{n}$ is a cube and for all measurable sets $E\subset Q$. In particular, there exists a constant $D$, depending only on $n$ and $[\omega]_{A_{2}}$, called the doubling constant for $\omega$ such that
@@
In particular, $L^{2}_{\omega}(\mathbb{R}^{n})$ is the Hilbert space of square-integrable functions on $\mathbb{R}^{n}$ with respect to $\mathrm{d}\omega$. We denote its norm by ${\lVert\cdot\rVert}_{2,\omega}$ and its inner product by $\langle\cdot,\cdot\rangle_{2,\omega}$. The class $\mathcal{D}(\mathbb{R}^{n})$ use of smooth ($C^{\infty}$), compactly supported test functions on $\mathbb{R}^{n}$, is dense in $L^{2}_{\omega}(\mathbb{R}^{n})$ as it is dense in $C_{c}(\mathbb{R}^{n})$, the space of continuous functions on $\mathbb{R}^{n}$ with compact support, and this latter space is dense in $L^{2}_{\omega}(\mathbb{R}^{n})$ as $\mathrm{d}\omega$ is a Radon measure on $\mathbb{R}^{n}$. Moreover, using the condition (2.1) with $p=2$, we have
@@ 2.3. The degenerate Laplacian
We define $-\Delta_{\omega}$ as the unbounded self-adjoint operator on $L^{2}_{\omega}(\mathbb{R}^{n})$ associated to the positive symmetric sesquilinear form on $H^{1}_{\omega}(\mathbb{R}^{n})\times H^{1}_{\omega}(\mathbb{R}^{n})$ defined by
@@
The operator $-\Delta_{\omega}$ is injective as the measure $\mathrm{d}\omega$ has infinite mass by (2.2). For all $\beta\in\mathbb{R}$, we define $(-\Delta_{\omega})^{\beta}$ as the self-adjoint operator $\mathbf{t^{\beta}}(-\Delta_{\omega})$, which is constructed via the functional calculus for sectorial operators, or equivalently, through the Borel functional calculus, since $-\Delta_{\omega}$ is self-adjoint. For further details, we refer to [RS80, McI86, Haa06]. Let the domain of $(-\Delta_{\omega})^{\beta}$ be denoted by $D((-\Delta_{\omega})^{\beta})$. Note that if $\beta>0$, then $D((-\Delta_{\omega})^{-\beta})$ is the range of $(-\Delta_{\omega})^{\beta}$ from its domain.
The operator $-\Delta_{\omega}$ generates a contractive holomorphic semi-group $(e^{z\Delta_{\omega}})_{z\in S_{\frac{\pi}{2}}}$ on the half-plane $S_{\frac{\pi}{2}}:=\left\{z\in\mathbb{C}\setminus\{0\}\,:|\mathrm{arg}(z)|<\frac{\pi}{2}\right\}=\left\{z\in\mathbb{C}\,:\mathrm{Re}(z)>0\right\}$. We now present the following lemma, which provides the gaussian upper and lower bounds for the positive-time heat kernels. These estimates are well known and we refer to [CUR14, AN25, Baa25] by setting $A=I_{n}$ in those references.
@@
Following the standard argument in [AT98, Lemma 19, p. 48] (more precisely its adaptation to the weighted setting in [CUR14, Theorem 4]), one derives the following Gaussian estimates for complex-time and time derivatives of the heat kernels.
@@
More properties hold but we point out that this is all we need for our proof of Theorem 1.1.
@@ 3.1. Key lemmas
Recall that the semigroup extends to a bounded operator on $L^{p}_{\omega}(\mathbb{R}^{n})$ for all $1\leq p\leq\infty$. We wish to obtain "$L^{p}-L^{q}$ bounds" when $p<q$. We record the following key lemma on the boundedness of the semigroup, with a bound involving a suitable power of $t$, provided the weight $\omega$ lies in a reverse Hölder class. We shall see that the open ended property of the reverse Hölder condition assumed in this proof will be used in an essential way.
@@
(3.1) $\left(\frac{1}{|Q|}\int_{Q}\omega^{\frac{q}{p}+\varepsilon}(x)\ \mathrm{d}x\right)^{\frac{1}{\frac{q}{p}+\varepsilon}}\leq C\frac{1}{|Q|}\int_{Q}\omega(x)\ \mathrm{d}x,$
(3.2) $\left(\frac{1}{|Q|}\int_{Q}\omega^{1-\frac{p^{\prime}}{q}\varepsilon}(x)\ \mathrm{d}x\right)^{\frac{1}{1-\frac{p^{\prime}}{q}\varepsilon}}\leq\frac{1}{|Q|}\int_{Q}\omega(x)\ \mathrm{d}x.$
We fix $\varepsilon\in(-\varepsilon_{\omega,p,q},\varepsilon_{\omega,p,q})$ and set $\gamma=\frac{1}{p}+\frac{\varepsilon}{q}\in(0,1)$. We fix $f\in L^{p}_{\omega^{1+(p/q)\varepsilon}}(\mathbb{R}^{n})\cap L^{2}_{\omega}(\mathbb{R}^{n})$ We fix $t>0$ and let $m\in\mathbb{Z}$ be the integer verifying $\frac{\sqrt{t}}{2}<\frac{1}{2^{m}}\leq\sqrt{t}$. Let $(Q_{k})_{k\in\mathbb{Z}^{n}}$ the family of dyadic cubes with side lengths $\frac{1}{2^{m}}$. Using the upper bound for the heat kernel in Lemma 2.3, we have
@@
Furthermore, we have for all $k,\ell\in\mathbb{Z}^{n}$,
@@
Step 1. We first show weak type bounds: namely, $\omega^{\frac{1}{p}+\frac{\varepsilon}{q}}\,T_{\delta,R}$ extends to a bounded operator from $L^{p}_{\omega^{1+(p/q)\varepsilon}}(\mathbb{R}^{n})$ to the Lorentz space $L^{q,\infty}_{dx}(\mathbb{R}^{n})$ with bounds independent of $\delta,R$ for all $1<p<q<\infty$ with $\alpha=\frac{n}{2}(\frac{1}{p}-\frac{1}{q})$ and $\varepsilon$ in a small interval $(-\tilde{\varepsilon}_{\omega,p,q},\tilde{\varepsilon}_{\omega,p,q})$ depending on $[\omega]_{RH_{\frac{q}{p}}}$, $p$, $q$ and $n$. For $\varepsilon=0$, this also holds when $p=1$.
@@
|$|\{\omega^{\frac{1}{p}+\frac{\varepsilon}{q}}\,|T_{\delta,R}f|>\lambda\}|$ |$\leq|\{\omega^{\frac{1}{p}+\frac{\varepsilon_{0}}{q_{0}}}\,|T_{\delta,a}f|>\frac{\lambda}{2}\}|+|\{\omega^{\frac{1}{p}+\frac{\varepsilon_{1}}{q_{1}}}\,|T_{a,R}f|>\frac{\lambda}{2}\}|$
@@
(3.6) |$|\{\omega^{\frac{1}{p}+\frac{\varepsilon}{q}}\,|T_{\delta,R}f|>\lambda\}|\leq 2C\,\lambda^{-q}.$
@@
(3.7) |$|\{\omega\,|T_{\delta,R}f|>\lambda\}|\leq 2C\,\lambda^{-q}.$
@@
Step 3. The third step consists in passing to the limit to obtain (1.1) and (1.2). We begin with the case $p>1$. The inequality (3.10) remains valid for all $f\in L^{2}_{\omega}(\mathbb{R}^{n})$ (as the inequality is obvious if $\|f\|_{L^{p}_{\omega}(\mathbb{R}^{n})}=\infty$). Now, we consider $f\in D((-\Delta_{\omega})^{\alpha})$. Using this, we have
@@
Step 4. To obtain the inequalities for $(-\Delta_{\omega})^{-\alpha}$, one can observe that $T_{\delta,R}f\to(-\Delta_{\omega})^{-\alpha}f$ in $L^{2}_{\omega}(\mathbb{R}^{n})$ when $f\in D((-\Delta_{\omega})^{-\alpha})$. Hence by the same arguments starting either from (3.10) or (3.7), we obtain the desired inequalities on $D((-\Delta_{\omega})^{-\alpha})\cap L^{p}_{\omega}(\mathbb{R}^{n})$. It remains to prove that, for all $1\leq p<\infty$, $D((-\Delta_{\omega})^{-\alpha})\cap L^{p}_{\omega}(\mathbb{R}^{n})$ is dense in $L^{p}_{\omega}(\mathbb{R}^{n})$. Since the domains form an increasing sequence of spaces in $-\alpha$, it suffices to prove it when $\alpha$ is an integer, say $N\in\mathbb{N}$, $N\geq 1$. We define $\Phi(z)=cz^{N}e^{-z}$, where $c>0$ is chosen so that

### v2

@@ 1. Introduction
Degenerate elliptic operators where the degeneracy is controlled by weights in the Muckenhoupt class have been introduced in [FKS82]. Since then, they have attracted attention with extension to parabolic equations [CS84b, CS84a, CS85, ISH99]. For $A_{2}(\mathbb{R}^{n})$ weights $\omega$, $\Delta_{\omega}$ can be defined as a maximal accretive operator on the weighted space $L^{2}_{\omega}(\mathbb{R}^{n})$. Much is known. For example, the associated semigroup is bounded on $L^{p}_{\omega}(\mathbb{R}^{n})$ for all $1\leq p\leq\infty$, it has $H^{\infty}$ functional calculus (equivalently, Borel functional calculus) if $1<p<\infty$. The particular study of its square root can be found in [CR15, CMR18] with a proof of the Kato problem on $L^{2}_{\omega}(\mathbb{R}^{n})$ and extension to some range of $p$, and even a "weighted theory" with weights in the Muckenhoupt classes for the measure w(x)dxw(x)\,dx.
@@
Inequalities (1.2) and (1.3) are a priori weighted Sobolev embeddings of a functional nature. Some more geometric weighted Sobolev-Gagliardo-Nirenberg inequalities were obtained in [LMP+10]. Namely, for all $1\leq p<q<\infty$ with $\frac{1}{2}=\frac{n}{2}(\frac{1}{p}-\frac{1}{q})$, $\|f\|_{L^{q}_{{\tilde{\omega}}^{q}}(\mathbb{R}^{n})}\lesssim\|\nabla f\|_{L^{p}_{\tilde{\omega}^{p}}(\mathbb{R}^{n})}$ when $\tilde{\omega}\in A_{p,q}(\mathbb{R}^{n})$ (which, again, requires $\tilde{\omega}^{p}\in RH_{\frac{q}{p}}(\mathbb{R}^{n})$).
We can only compare these results with ours when $\alpha=1/2$ in the range of $p$ for which $\|\nabla f\|_{L^{p}_{\omega}(\mathbb{R}^{n})}$ dominates or is dominated by $\|(-\Delta_{\omega})^{1/2}f\|_{L^{p}_{\omega}(\mathbb{R}^{n})}$. When $p=2$, one has $\|(-\Delta_{\omega})^{1/2}f\|_{L^{2}_{\omega}(\mathbb{R}^{n})}=\|\nabla f\|_{L^{2}_{\omega}(\mathbb{R}^{n})}$ by construction. Moreover, deducing our results from these or other known results (e.g., [MW74]), or conversely, requires additional restrictions on the class of weights and on the dimension, as we shall explain.
As a matter of fact, we will show that our reverse Hölder assumption is sharp in the sense that it is implied by (1.2). At the level of the semigroup, we shall see that with the assumptions of the theorem, then $\omega\in RH_{\frac{q}{p}}(\mathbb{R}^{n})$ is equivalent to $e^{t\Delta_{w}}:L^{p}_{\omega}(\mathbb{R}^{n})\to L^{q}_{{\omega}^{q/p}}(\mathbb{R}^{n})$ with bound on the order $t^{-\alpha}$ for all $t>0$. This polynomial bound is the key point of our method.
Our method also allows for various generalizations. In [O’N63], O’Neil showed that the classical unweighted $L^{p}-L^{q}$ estimates for Riesz potentials (that is, the fractional Laplacians of negative order) can be improved to $L^{p}-L^{q,p}$ estimates, where $L^{q,p}$ is a Lorentz space, using the translation invariance of the operators. Later, it was observed by Tartar (see [TAR98]) that this is a consequence of real interpolation more than translation invariance. Here too, real interpolation applies and yields two apparently distinct improvements for the weighted fractional Laplacian, see Theorem 3.15 and Theorem 3.17, in agreement with observations made in [CMM22]. In particular, we shall show that $w^{\frac{1}{p}}(-\Delta_{\omega})^{-\alpha}$ is bounded from $L^{p}_{\omega}(\mathbb{R}^{n})$ to the unweigthed Lorentz space $L^{q,p}(\mathbb{R}^{n})$ under the same assumption as for (1.2).
@@
As a motivation for this work, bounds for fractional weighted Laplacians are used to construct solutions to second-order degenerate parabolic equations with unbounded lower-order terms [BAA25]. They are also used to establish a boundedness result for degenerate parabolic Riesz transforms. The second author will address this in a forthcoming work.
@@
Throughout this paper, we adopt the following notation.
  1. $\blacklozenge$
    We fix an integer $n\geq 1$.
  2. $\blacklozenge$
    We use the notation $\mathrm{d}x$ for the Lebesgue measure on $\mathbb{R}^{n}$.
  3. $\blacklozenge$
    For any Lebesgue measurable set $E\subset\mathbb{R}^{n}$, we denote its Lebesgue measure by $|E|$.
  4. $\blacklozenge$
    By convention, the notation $C(a,b,\ldots)$ denotes a constant depending only on the parameters $(a,b,\ldots)$.
  5. $\blacklozenge$
    All cubes $Q\subset\mathbb{R}^{n}$ considered in this paper are assumed to have sides parallel to the coordinate axes.
@@ Acknowledgements
We want to thank Chema Martell for bringing our attention to [CMM22] concerning Lorentz spaces after we completed this work and Carlos Pérez for enlightening discussions. We also thank the anonymous referee for providing suggestions that enhanced the presentation.
@@ 2.1. Muckenhoupt weights
(2.1) $\omega]_{A_{p}}:=\sup_{Q\subset\mathbb{R}^{n}}\left(\frac{1}{|Q|}\int_{Q}\omega(x)\ \mathrm{d}x\right)\left(\frac{1}{|Q|}\int_{Q}\omega^{1-p^{\prime}}(x)\ \mathrm{d}x\right)^{p-1}<\infty,$
@@
$\left(\frac{1}{|Q|}\int_{Q}\omega^{q}(x)\ \mathrm{d}x\right)^{\frac{1}{q}}\leq\frac{C}{|Q|}\int_{Q}\omega(x)\ \mathrm{d}x,$
for some constant $C<\infty$ and for all cubes $Q\subset\mathbb{R}^{n}$. The infimum of such constants $C$ is what we denote by $\omega]_{RH_{q}}$. For $q=\infty$, the reverse Hölder class $RH_{\infty}(\mathbb{R}^{n})$ is defined as the set of all weights $\omega$ verifying
$\omega(y)\leq\frac{C}{|Q|}\int_{Q}\omega(x)\ \mathrm{d}x,\quad\text{for almost every}\ y\in Q,$
@@
$\omega]_{A_{p,q}}:=\sup_{Q\subset\mathbb{R}^{n}}\left(\frac{1}{|Q|}\int_{Q}\omega^{q}(x)\ \mathrm{d}x\right)^{\frac{1}{q}}\left(\frac{1}{|Q|}\int_{Q}\omega^{-p^{\prime}}(x)\ \mathrm{d}x\right)^{\frac{1}{p^{\prime}}}<\infty,$
@@
It is well known (see, for instance, [GRA08, Example 7.1.7]) that a weight of the form $\omega(x)=|x|^{\beta}$ belongs to the Muckenhoupt class $A_{p}(\mathbb{R}^{n})$ for $p>1$ if and only if $n<\beta<n(p-1)$. Furthermore, according to point (6) of Proposition 2.2 above, a weight $\omega$ belongs to the intersection $A_{2}(\mathbb{R}^{n})\cap RH_{\frac{q}{p}}(\mathbb{R}^{n})$ for $p<q$ if and only if $\omega^{\frac{q}{p}}\in A_{1+\frac{q}{p}}(\mathbb{R}^{n})$. As a result, a typical example of a weight $\omega$ considered in the statements of the results in this paper, satisfying $\omega\in A_{2}(\mathbb{R}^{n})\cap RH_{\frac{q}{p}}(\mathbb{R}^{n})$ with $p<q$, is given by
@@ 2.2. Function spaces
Throughout this paper, $\omega$ denotes a fixed weight belonging to the Muckenhoupt class $A_{2}(\mathbb{R}^{n})$. By definition (2.1) with $p=2$, the weight $\omega^{-1}$ is also in the Muckenhoupt class $A_{2}(\mathbb{R}^{n})$ with $\omega^{-1}]_{A_{2}}=[\omega]_{A_{2}}$. We introduce the measure $\mathrm{d}\omega:=\omega(x)\mathrm{d}x$ and if $E\subset\mathbb{R}^{n}$ a Lebesgue measurable set, we write $\omega(E)$ instead of $\int_{E}\mathrm{d}\omega$. For all $x\in\mathbb{R}^{n}$ and $r>0$, we set $\omega_{r}(x):=\omega(B(x,\sqrt{r}))$.
It follows from (2.1) with $p=2$ that there exits constants $\eta\in(0,1)$ and $\beta>0$, depending only on $n$ and $\omega]_{A_{2}}$, such that
@@
whenever $Q\subset\mathbb{R}^{n}$ is a cube and for all measurable sets $E\subset Q$. In particular, there exists a constant $D$, depending only on $n$ and $\omega]_{A_{2}}$, called the doubling constant for $\omega$ such that
@@
In particular, $L^{2}_{\omega}(\mathbb{R}^{n})$ is the Hilbert space of square-integrable functions on $\mathbb{R}^{n}$ with respect to $\mathrm{d}\omega$. We denote its norm by ${\lVert\cdot\rVert}_{2,\omega}$ and its inner product by $\langle\cdot,\cdot\rangle_{2,\omega}$. The class $\mathcal{D}(\mathbb{R}^{n})$ consists of smooth ($C^{\infty}$), compactly supported test functions on $\mathbb{R}^{n}$, is dense in $L^{2}_{\omega}(\mathbb{R}^{n})$ as it is dense in $C_{c}(\mathbb{R}^{n})$, the space of continuous functions on $\mathbb{R}^{n}$ with compact support, and this latter space is dense in $L^{2}_{\omega}(\mathbb{R}^{n})$ as $\mathrm{d}\omega$ is a Radon measure on $\mathbb{R}^{n}$. Moreover, using the condition (2.1) with $p=2$, we have
@@ 2.3. The degenerate Laplacian
We define $\Delta_{\omega}$ as the unbounded self-adjoint operator on $L^{2}_{\omega}(\mathbb{R}^{n})$ associated to the positive symmetric sesquilinear form on $H^{1}_{\omega}(\mathbb{R}^{n})\times H^{1}_{\omega}(\mathbb{R}^{n})$ defined by
@@
The operator $\Delta_{\omega}$ is injective as the measure $\mathrm{d}\omega$ has infinite mass by (2.4). For all $\beta\in\mathbb{R}$, we define $(-\Delta_{\omega})^{\beta}$ as the self-adjoint operator $\mathbf{t^{\beta}}(-\Delta_{\omega})$, that is, the operator obtained by applying the function $f_{\beta}(z)=z^{\beta}$ to $\Delta_{\omega}$. This construction can be made either via the functional calculus for sectorial operators or, equivalently, via the Borel functional calculus, since $\Delta_{\omega}$ is self-adjoint. For further details, we refer to [RS80, MCI86, HAA06]. Let the domain of $(-\Delta_{\omega})^{\beta}$ be denoted by $D((-\Delta_{\omega})^{\beta})$. Note that if $\beta>0$, then $D((-\Delta_{\omega})^{-\beta})$ is the range of $(-\Delta_{\omega})^{\beta}$ from its domain.
The operator $\Delta_{\omega}$ generates a contractive holomorphic semi-group $(e^{z\Delta_{\omega}})_{z\in S_{\frac{\pi}{2}}}$ on the half-plane $S_{\frac{\pi}{2}}:=\left\{z\in\mathbb{C}\setminus\{0\}\,:|\mathrm{arg}(z)|<\frac{\pi}{2}\right\}=\left\{z\in\mathbb{C}\,:\mathrm{Re}(z)>0\right\}$. We now present the following lemma, which provides the Gaussian upper and lower bounds for the positive-time heat kernels. These estimates are well known and we refer to [CR14, AN25, BAA26] by setting $A=I_{n}$ in those references.
@@
Following the standard argument in [AT98, Lemma 19, p. 48] (more precisely its adaptation to the weighted setting in [CR14, Theorem 4]), one derives the following Gaussian estimates for complex-time and time derivatives of the heat kernels.
@@
More properties, such as the boundedness of the $H^{\infty}$-calculus of $\Delta_{\omega}$ and of its Riesz transform on $L^{p}_{\omega}(\mathbb{R}^{n})$ for $p\neq 2$, also hold [CMR18]. However, we point out that the above results are all we need for the proof of Theorem 1.1.
@@ 3.1. Key lemmas
Recall that the semigroup extends to a bounded operator on $L^{p}_{\omega}(\mathbb{R}^{n})$ for all $1\leq p\leq\infty$. We wish to obtain $L^{p}-L^{q}$ bounds" when $p<q$. We record the following key lemma on the boundedness of the semigroup, with a bound involving a suitable power of $t$, provided the weight $\omega$ lies in a reverse Hölder class. We shall see that the open ended property of the reverse Hölder condition assumed in this proof will be used in an essential way.
@@
(3.2) $\left(\frac{1}{|Q|}\int_{Q}\omega^{\frac{q}{p}+\varepsilon}(x)\ \mathrm{d}x\right)^{\frac{1}{\frac{q}{p}+\varepsilon}}\leq\frac{C}{|Q|}\int_{Q}\omega(x)\ \mathrm{d}x,$
(3.3) $\left(\frac{1}{|Q|}\int_{Q}\omega^{1-\frac{p^{\prime}}{q}\varepsilon}(x)\ \mathrm{d}x\right)^{\frac{1}{1-\frac{p^{\prime}}{q}\varepsilon}}\leq\frac{C}{|Q|}\int_{Q}\omega(x)\ \mathrm{d}x.$
We fix $\varepsilon\in(-\varepsilon_{\omega,p,q},\varepsilon_{\omega,p,q})$ and set $\gamma=\frac{1}{p}+\frac{\varepsilon}{q}\in(0,1)$. We fix $f\in L^{p}_{\omega^{1+(p/q)\varepsilon}}(\mathbb{R}^{n})\cap L^{2}_{\omega}(\mathbb{R}^{n})$ . We fix $t>0$ and let $m\in\mathbb{Z}$ be the integer verifying $\frac{\sqrt{t}}{2}<\frac{1}{2^{m}}\leq\sqrt{t}$. Let $(Q_{k})_{k\in\mathbb{Z}^{n}}$ the family of dyadic cubes with side lengths $\frac{1}{2^{m}}$. Using the upper bound for the heat kernel in Lemma 2.6, we have
@@
Furthermore, for all $k,\ell\in\mathbb{Z}^{n}$ with $k\neq\ell$, we have
@@
Step 1. We first show weak type bounds: namely, $\omega^{\frac{1}{p}+\frac{\varepsilon}{q}}\,T_{\delta,R}$ extends to a bounded operator from $L^{p}_{\omega^{1+(p/q)\varepsilon}}(\mathbb{R}^{n})$ to the Lorentz space $L^{q,\infty}_{dx}(\mathbb{R}^{n})$ with bounds independent of $\delta,R$ for all $1<p<q<\infty$ with $\alpha=\frac{n}{2}(\frac{1}{p}-\frac{1}{q})$ and $\varepsilon$ in a small interval $(-\tilde{\varepsilon}_{\omega,p,q},\tilde{\varepsilon}_{\omega,p,q})$ depending on $\omega]_{RH_{\frac{q}{p}}}$, $p$, $q$ and $n$. For $\varepsilon=0$, this also holds when $p=1$.
@@
$|\{\omega^{\frac{1}{p}+\frac{\varepsilon}{q}}\,|T_{\delta,R}f|>\lambda\}|$ $\leq|\{\omega^{\frac{1}{p}+\frac{\varepsilon_{0}}{q_{0}}}\,|T_{\delta,a}f|>\frac{\lambda}{2}\}|+|\{\omega^{\frac{1}{p}+\frac{\varepsilon_{1}}{q_{1}}}\,|T_{a,R}f|>\frac{\lambda}{2}\}|$
@@
(3.8) $|\{\omega^{\frac{1}{p}+\frac{\varepsilon}{q}}\,|T_{\delta,R}f|>\lambda\}|\leq 2C\,\lambda^{-q}.$
@@
(3.9) $|\{\omega\,|T_{\delta,R}f|>\lambda\}|\leq 2C\,\lambda^{-q}.$
@@
Step 3. The third step consists in passing to the limit to obtain (1.2) and (1.3). We begin with the case $p>1$. The inequality (3.12) remains valid for all $f\in L^{2}_{\omega}(\mathbb{R}^{n})$ (as the inequality is obvious if $\|f\|_{L^{p}_{\omega}(\mathbb{R}^{n})}=\infty$). Hence, substituting $f$ by $(-\Delta_{\omega})^{\alpha}f\in L^{2}_{\omega}(\mathbb{R}^{n})$ for $f\in D((-\Delta_{\omega})^{\alpha})$, we have
@@
Step 4. To obtain the inequalities for $(-\Delta_{\omega})^{-\alpha}$, one can observe that $T_{\delta,R}f\to(-\Delta_{\omega})^{-\alpha}f$ in $L^{2}_{\omega}(\mathbb{R}^{n})$ when $f\in D((-\Delta_{\omega})^{-\alpha})$. Hence by the same arguments starting either from (3.12) or (3.9), we obtain the desired inequalities on $D((-\Delta_{\omega})^{-\alpha})\cap L^{p}_{\omega}(\mathbb{R}^{n})$. It remains to prove that, for all $1\leq p<\infty$, $D((-\Delta_{\omega})^{-\alpha})\cap L^{p}_{\omega}(\mathbb{R}^{n})$ is dense in $L^{p}_{\omega}(\mathbb{R}^{n})$. Since the domains form an increasing sequence of spaces in $\alpha$, it suffices to prove it when $\alpha$ is an integer, say $N\in\mathbb{N}$, $N\geq 1$. We define $\Phi(z)=cz^{N}e^{-z}$, where $c>0$ is chosen so that